Given that f(x)=3x+1, g(x)=x^2-2x-6 and h(x)=x^3, find each:
(f o g)(1/3)

Answer by drk(1908)

First, we find g(1/3) as
g(1/3) = (1/3)^2 - 2(1/3) - 6 = 1/9 - 2/3 - 6 = 5 5/9 = 50/9
second
(f of g)(1/3) = 3(50/9) + 1 = 50/3 + 1 = 17 2/3